OpenStudy (anonymous):

Find the exact value of the expression. Do not use a calculator. If tan θ = 9, find the exact value of cot \[[\frac{ \pi }{ 2 } - \theta]\] . Please explain how to solve

OpenStudy (anonymous):

from the definition of cotangent... \[\cot \left[\frac{\pi}{2}-\theta\right]=\tan \theta\] so if \(\tan \theta = 9\) then \[\cot \left[\frac{\pi}{2}-\theta\right]=9\]